OpenStudy (anonymous):

John began his job making $25 the first day. After that he was paid $6.75 per hour. Write the linear equation for this problem in y=mx+b form

OpenStudy (anonymous):

This is hard to solve for each variable, but I think all they want you to do is answer: \[y = \frac{675}{100}\left( x \right) + 25 \rightarrow y = \frac{25}{4}\left( x \right) + 25\] The equation on the right being simplified, where y = amount earnt and x = hours worked.

OpenStudy (anonymous):

thank you for the help i actually looked at my note and figured it out but thankss caue now i know i am doing it right :)

OpenStudy (anonymous):

Sorry it took so long, i was unsure too :) cool, thanks

OpenStudy (anonymous):

its no problem and :)